If the kinetic energy of an electron is given as “K” Joules with an uncertainty of 6*10^-6, then calculate the expression of (1)uncertainty in velocity and the (2)uncertainty in position.Express your answer in terms of “K” joules. 

Grade:11

1 Answers

Vikas TU
7 years ago
From uncertainity condition:
 
dp * dx > = h/4pi 
dp and dx are the uncertainities in momentum and position.
 
  1.  
    uncertainity invelocity=>
    given kinetic enrgy as: K = 1/2mv^2 = pv/2
    p = 2K/v
    put in uncertainity condition,
    we get,
    d(2K/v) * dx = h/4pi